# A systematic comparison of measures for k-anonymity in networks

### v1

@@ Abstract
Privacy-aware sharing of network data is a difficult task due to the interconnectedness of individuals in networks. An important part of this problem is the inherently difficult question of how in a particular situation the privacy of an individual node should be measured. To that end, in this paper we propose a set of aspects that one should consider when choosing a measure for privacy. These aspects include the type of desired privacy and attacker scenario against which the measure protects, utility of the data, the type of desired output, and the computational complexity of the chosen measure. Based on these aspects, we provide a systematic overview of existing approaches in the literature. We then focus on a set of measures that ultimately enables our objective: sharing the anonymized full network dataset with limited disclosure risk. The considered measures, each based on the concept of $k$-anonymity, account for the structure of the surroundings of a certain node and differ in completeness and ‘reach’ of the structural information taken into account. We present a comprehensive theoretical characterization as well as comparative empirical experiments on a wide range of real-world network datasets with up to millions of edges. We find that the choice of the measure has an enormous effect on aforementioned aspects. Most interestingly, we find that the most effective measures consider a greater node vicinity, yet utilize minimal structural information and thus use minimal computational resources. This finding has important implications for researchers and practitioners, who may, based on the recommendations given in this paper, make an informed choice on how to safely share large-scale network data in a privacy-aware manner.
Keywords Complex networks  $\cdot$ Privacy  $\cdot$ $k$-Anonymity
@@ 1 Introduction
Network science is a research field centered around modelling real-world phenomena using nodes, representing entities, and edges representing some type of meaningful connections between these entities. Various methods have been introduced to study networks, for example to find important nodes [1], communities [2] or anomalies [3]. These methods can be applied to model disease spread [4], measure segregation [5, 6], or identify suspicious entities in, for example, social networks [7]. For such research purposes it is beneficial to have access to networks that contain realistic information on individual entities, for example about people. While often social media platforms such as Meta/Facebook or X/Twitter are used, recently there has been a rise in attention for population-scale network analysis where the set of nodes represents the entire population of a country [8, 9]. Sharing such large-scale social networks of individuals for research purposes comes with considerable privacy risks, as individuals might be identified and sensitive data may be derived. Even after removing identifiers, i.e., pseudonymization, a network is not sufficiently protected, as the structure of the network itself can be used to identify individuals [10, 11, 12]. Subsequently, social network data is often not shared or can be accessed by a limited number of researchers under strict conditions. This to some extent hinders the transition to open science.
@@
The problem we focus on in this work is that of sharing data while preserving privacy. This has many aspects that should be taken into account when choosing an approach to address this problem. This includes 1) the type of desired output, 2) the ‘attacker scenarios’ against which should be protected, and 3) the utility of the data in terms of which data properties should be preserved and how the data can be used. The most commonly used methods for the protection of network data have initially been introduced in the context of Statistical Disclosure Control (SDC) [13, 14] aiming to protect tabular data, i.e., data where each entity is described by a number of attributes. In this field, one distinguishes between quasi-identifiers, attributes which may be known to the attacker, who aims to de-anonymize the network, and could be used for identification, and sensitive attributes, which are unknown and the attacker aims to obtain.
@@
The most commonly used approaches in SDC are differential privacy [15, 16] and $k$-anonymity [17]. In differential privacy, a mechanism gives noisy answers to user queries such that the privacy of entities is preserved. For $k$-anonymity, the aim is to create a version of the data such that for each entity there are at least $k$ equivalent candidates, so that no single individual can be identified. Typically a value of $k>1$ is used. This is achieved by generalizing quasi-identifiers, or suppressing sensitive values in the dataset. However, in a $k$-anonymous dataset it would still be possible for an attacker to obtain information if all $k$ entities have the same sensitive attribute(s) or if the distribution in a certain class deviates from the distribution over the entire dataset. This is counteracted by extending $k$-anonymity with $\ell$-diversity [18], $t$-closeness [19] or ($\alpha$, $k$)-anonymity [20]. Other approaches are incorporating synthetic data [21] by either generating a new dataset from scratch or combining synthetic data with the actual data. Another approach to address this problem is anatomy [22], in which the quasi-identifiers and sensitive attributes are shared in separate tables such that it can not be inferred which set of sensitive attributes correspond to which set of quasi-identifiers.
SDC methods originally designed for tabular data have been adjusted to the case of network data. Several generically applicable approaches based on differential privacy [23, 24, 25] as well as $k$-anonymity [26, 27, 11, 28] have been introduced. Other approaches specific to networks are randomization [29, 30, 31] and clustering [32, 27, 33, 34, 35, 36, 37]. As we will see in Section 3, we can further categorize these approaches based on the type of output they generate.
The approach we focus on in this paper is $k$-anonymity. This approach ultimately enables one to share an anonymized version of the full network data while protecting against reidentification risk, attaining the goal of sharing an anonymized version of the network data. When using $k$-anonymity to this end, one needs to chose both an anonymization algorithm and a definition for equivalence, which we refer to as the anonymity measure. The goal of the anonymization algorithm is to achieve a $k$-anonymous network by means of perturbation, the output of which can safely be shared. In this network, each node is equivalent to at least $k-1$ other nodes, given the chosen definition for equivalence. While the anonymization step is an important one, in this paper we specifically focus on the measurement of $k$-anonymity.
Various anonymity measures can be used for node equivalence. These measures differ in the completeness and reach of the structural information they take into account. As a result, they differ in their ‘strictness’ of when two nodes are equivalent. Existing measures are typically based on the degree of a node [26] or nearby nodes [38], the ego-network structure [11, 12] or the exact structural position of a node in the network [28, 39]. Each measure essentially corresponds to an attacker scenario in which we assume the attacker has a certain amount and type of information on an entity. Stricter measures would protect against more attacker scenarios, but at the same time they add stronger requirements for a network to be anonymous. While the anonymization method can be chosen separately from the anonymization measure, in general it is more difficult to obtain an anonymous network for a stricter measure by means of perturbation. As more perturbations might be required to obtain an anonymized network, choosing a more strict measure may in turn result in lower data utility. On the other hand, if the network is protected against a lenient attacker scenario, one takes the risk of underestimating the attacker knowledge and the attacker may still be able to identify entities in the anonymized network.
It is clear that the choice of anonymity measure is of high importance. While many existing works aim to achieve $k$-anonymous network for various measures [26, 27, 11, 28], there is not yet a systematic comparison of these measures and how the choice of measure impacts the attained privacy of a $k$-anonymous network.
@@
As a key part of this paper, we compare different measures for $k$-anonymity and aim to understand them both theoretically and empirically by comparing them on a wide range of empirical networks. We distinguish between these measures based on how far they ‘reach’, i.e., up to where they take structural information into account, and how complete the structural information is. When we empirically compare the measures, we find that measures with a larger reach, even with minimal structural information, lead to higher risk of disclosure than having perfect structural knowledge over the direct neighborhood of a node.
@@
  • We provide an overview of aspects that should be taken into account when sharing network data while preserving privacy of entities represented (Section 2).
  • We give a categorization of state-of-the-art approaches in the literature (Section 3).
@@
  • We summarize existing measures for $k$-anonymity in networks (Section 4).
@@
  • We theoretically compare the measures based on reach and the completeness of the information they take into account, creating a formal strictness-based ordering of measures (Section 5).
@@
  • We empirically compare these measures in terms of both anonymity observed and runtime on a wide range of real-world network datasets (Section 6).
The remainder of the paper is structured as follows. In Section 2, we summarize the aspects that should be taken into account when choosing an approach and measure for sharing network data while preserving the privacy of entities represented in it. Next, Section 3 gives an overview and categorization of approaches from the literature. Then, in Section 4 we give an overview of the measures for $k$-anonymity introduced in the literature. We compare a set of representative measures both theoretically and empirically on a wide range of networks in Sections 5 and 6. Lastly, we conclude the paper and indicate possible future work in Section 7.
2 Aspects of sharing sensitive network data
When deciding which approach to use for sharing network data while preserving privacy of entities represented, there are three important aspects to keep in mind: utility, privacy and computation time. We describe all three of these aspects in detail below.
2.1 Utility
For the utility of the data, we distinguish between two aspects. First, how the data can be shared in a useful manner, i.e., what is the desired type of output, and second which structural network properties should be preserved.
Starting with the first, we can distinguish between four types of output: 1) an interactive mechanism that allows users to ask queries about the network, 2) an intermediate representation of the network from which new networks can be generated, 3) a synthetic network or 4) a perturbed version of the network which can immediately be analyzed. While the first type is interactive, the other are non-interactive [15] as they share a version of the data once and do not require further interaction with the user.
The choice of output relates to the trade-off between data utility and privacy. When using an interactive approach where a user is allowed to submit queries, the information shared is tractable. Additionally, one could choose to not implement queries which may be a threat for privacy. A non-interactive approach would allow the user to run any query. However, this gives an advantage to possible attackers. While methods usually assume a certain attacker scenario, an attacker could come up with an unforeseen method, e.g., twin nodes as proposed in [12], or gather more data than foreseen in the attacker scenario, e.g., about centrality [40], which can be used to extract sensitive information. Moreover, when a non-interactive approach is used, this allows users to get more data from the network as they are not restricted to the queries that are implemented for the interactive mechanism.
A second utility aspect one needs to consider is which properties of the network are important to preserve. These properties could be global, concerning properties considering the graph as a whole such as the diameter; local, concerning properties surrounding the nodes such as clustering and degree; or at the meso level, for example accounting for the community structure. With this in mind, one could choose an approach that better preserves these properties as some methods may by their nature be more likely to destroy or preserve certain properties.
2.2 Privacy
For the privacy aspect, one needs to determine four things: 1) what the sensitive information in the network is, 2) what a realistic amount of structural attacker knowledge is 3) if there are any properties of the network that can be used by the attacker and 4) if the attacker can play a role in the network to cause alterations. We discuss each in detail below.
2.2.1 Sensitive information
Overall, works distinguish between node privacy and edge privacy [41, 42]. When protecting for node privacy, an attacker should not be able to infer the identity of a node. For edge privacy, the attacker should not be able to infer if there exist a relationship between two entities. Edge privacy would not imply node privacy, but node privacy, which is the focus for this paper, conditionally implies edge privacy. Intuitively, when the identities of the two nodes an edge connects are unknown, the existence of a link can not be inferred, hence assuring edge privacy. However, edge privacy is not guaranteed if the existence of the connection can be logically inferred, e.g., if for all pairs of candidate nodes there exists an edge. These notions can be linked to identity and attribute disclosure in SDC, as briefly covered in Section 1. Node privacy would prevent an entity from being identified in the dataset and thus prevent identity disclosure, while edge privacy would prevent one from learning about connections between entities, which can be seen as an attribute disclosure.
2.2.2 Attacker knowledge
To estimate the attacker scenario, we need to assess which type of knowledge the attacker can obtain, in which amount, and which properties they can use to obtain more sensitive information. In the context of $k$-anonymity, which we discuss in detail in Section 4, the choice of attacker scenario plays an important role in choosing a measure. Examples are the degree-based attack [26, 43, 44], where the attacker has knowledge of degrees of nodes, or a neighborhood attack [11, 45], where we assume a possible attacker has knowledge of a nodes’ surrounding neighborhood. Considering the amount of attacker knowledge, assuming more knowledge overall protects against more attacker scenarios, resulting in a stronger privacy guarantee. However, when one assumes an attacker has more knowledge, it may be more difficult to achieve anonymity as in an anonymization approach based on perturbation more changes might be required, which in turn could result in lower data utility. By choosing a less strict attacker scenario, anonymity may be easier to achieve, which may indirectly result in higher utility. However, it may not ensure sufficient privacy. Hence, it is very important to take into account a realistic attacker scenario.
2.2.3 Types of attacker scenarios
Various types of attacker scenarios have been introduced in literature [46, 47]. These works distinguish between seed based and seedless attacks. For seed based attacks, it is assumed that the attacker knows a number of user identities which are referred to as seeds. Most approaches include a propagation or cascade step which uses the identified entities to de-anonymize more entities [48, 49]. These works assume an attacker has access to different social networks partially overlapping in people present. These networks can be used to link similar nodes to each other. If this can be done properly, this could be a threat to privacy. However, it has been shown that this approach needs many high quality seed nodes [49], which may not always be realistic. Seedless attacks assume that an attacker does not need any seeds to start. Instead, structural network properties are used for identification.
2.2.4 Attacker role
Lastly, literature distinguishes between active and passive attacks [41, 50, 47]. For active attacks, also referred to as sybil attacks, the assumption is that an attacker can make changes to the network, e.g., by making friends on a social media platform. After an anonymized version of the network is obtained, the attacker can use these sybil nodes to identify themselves, and possibly cascade this information to de-anonymize more nodes. When this is possible, this is an important aspect to take into account in the attacker scenario. For passive attacks, which is the type of attack assumed in this paper, either a single network is shared, or it is assumed an attacker is not able to alter the network to such a large extend, e.g., for an electric grid network.
2.3 Computation time
When choosing an approach, one should also take into account whether it is computationally feasible to use this on the network data at hand. Some anonymization algorithms are more computationally expensive than others, making them possibly unsuitable for larger networks. For example, the problem of achieving $k$-anonymity with as few changes as possible is proven to be NP-hard when using 1-neighborhood isomorphism as measure [51], hence an approximation algorithm has to be considered. When using the degree as measure, determining the smallest set of alterations required to anonymize the graph by means of perturbation can be done efficiently. For example, the algorithm introduced in [26] runs in quadratic time in terms of the number of nodes. There essentially exists a trade-off between privacy and computational cost which should be taken into account when choosing an approach for a given network dataset.
3 Methods for privacy aware sharing of networks
In this section, we summarize approaches introduced in literature for achieving privacy aware sharing of networks. We categorize the approaches based on the types of output mentioned in Section 2.1 and discuss each approach accordingly. The categories and corresponding properties can be found in Table 1.
@@ 3 Methods for privacy aware sharing of networks
Differential privacy: queries [24, 52, 53, 42] $\checkmark$ $\lozenge$ $\checkmark$
@@
Clustering [32, 33, 34, 35] $\checkmark$ $\checkmark$
@@
Table 1: Categorization of methods for privacy-aware sharing of networks. $\checkmark$ indicates the method is in this category, $\lozenge$ that it could be in this category, but is less commonly studied as such.
3.1 Interactive
The first category that we discuss is that of the interactive approaches, which primarily consists of differential privacy based approaches. These approaches allow users to ask queries about the network dataset [56]. The user receives an answer to which noise is added, such that it preserves either node or edge privacy. The attacker scenario assumed is that the attacker has knowledge about the entire network except for one node or edge. Based on the answer, it should not be possible to infer whether a specific edge, when protecting edge privacy, or node, when protecting node privacy, exists in the dataset. The amount of noise added is based on the sensitivity of the query and ensures the privacy is guaranteed. Most works on differential privacy focus on edge privacy [23, 25, 54] and some on node privacy [53, 57]. As the (non-)existence of a node can potentially have more effect on the query than the (non-)existence of an edge, it is overall more difficult to ensure node privacy. Hence, more noise should be added to achieve node privacy.
When applying differential privacy to networks, there are two approaches, each providing different types of output. One aims to answer a specific user query with as little noise as possible [24, 52, 53, 42]. In this case, the approach is interactive and users are dependent on the queries that are implemented. A disadvantage of this approach is that when more queries are posed, more noise should be added to ensure differential privacy [42]. This strongly reduces the quality of the answers and with that the utility of the resulting dataset.
The other set of differential privacy approaches aims to create a synthetic network based on the output of a query [23, 25, 54]. This can for example be done by generating a synthetic dK-graph [58] (see also Section 3.3) based on a differentially private joint degree distribution of the network. This distribution can be obtained from the original graph by determining for each edge the degree of the nodes it connects, and counting the occurrence of each such pair. For this type of approach, the quality of the resulting network depends both on the quality of the answer provided by the differential privacy mechanism and the graph model used, in this case the dK-graph.

### v2

# A systematic comparison of measures for publishing $k$-anonymous social network data

@@ Abstract
Sharing or publishing social network data while accounting for privacy of individuals is a difficult task due to the interconnectedness of nodes in networks. A key question in $k$-anonymity, a widely studied notion of privacy, is how to measure the anonymity of an individual, as this determines the attacker scenarios one protects against. In this paper, we systematically compare the most prominent anonymity measures from the literature in terms of the ‘completeness’ and ‘reach’ of the structural information they take into account. We present a theoretical characterization and a distance-parametrized strictness ordering of the existing measures for $k$-anonymity in networks. In addition, we conduct empirical experiments on a wide range of real-world network datasets with up to millions of edges. Our findings reveal that the choice of the measure significantly impacts the measured level of anonymity and hence the effectiveness of the corresponding attacker scenario, the privacy vs. utility trade-off, and computational cost. Surprisingly, we find that the anonymity measure representing the most effective attacker scenario considers a greater node vicinity yet utilizes only limited structural information and therewith minimal computational resources. Overall, the insights provided in this work offer researchers and practitioners practical guidance for selecting appropriate anonymity measures when sharing or publishing social network data under privacy constraints.
keywords:
$k$-Anonymity, Social Networks, Privacy, Algorithms, Utility, Statistical Disclosure Control
@@ 1 Introduction
Network science is a research field centered around modeling real-world phenomena using nodes, representing entities, and edges representing some type of meaningful connection between these entities. Various methods have been introduced to study networks, for example, to find important nodes [1], communities [2] or anomalies [3]. In social networks, these methods can be applied to model disease spread [4], measure segregation [5, 6], or identify suspicious entities in (social) networks [7]. For such research purposes it is beneficial to have access to networks that contain realistic information on individual entities, for example, about people. Often data from social media platforms such as Meta/Facebook or X/Twitter are used to construct networks, but also population-scale networks generated from official government register data where the set of nodes represents the entire population of a country [8, 9] are frequent sources for social network analysis research.
What abovementioned network datasets have in common is that sharing or publicly publishing such large-scale social networks, even if it is for research purposes, comes with considerable privacy risks. Even after personally identifiable information is removed and replaced with artificial identifiers, i.e., pseudonymization, a network is not sufficiently protected against disclosure risk, as the structure of the network itself can be used to identify individuals [10, 11, 12]. As a consequence, real-world social network data is often not shared or can only be accessed under strict conditions. At the same time, promoted by, e.g., the FAIR data initiative [13], researchers are often expected to digitally publish social network datasets used together with their paper for reproducibility and reusability [14]. It is important to note that in this concrete scenario, it is not possible to adopt techniques based on differential privacy. Such methods require generating synthetic data at the cost of analytical fidelity, or only work in “online” systems that hold the original data and dynamically add noise as the graph is queried [15, 16]. Concretely, in this work we ultimately focus on obtaining a publishable file of a network dataset that remains representative of the real system, while protecting individual privacy.
@@
The problem of sharing social network data in some form while accounting for the privacy of people represented has four important aspects that should be taken into account. This includes 1) the type of desired output, 2) the ‘attacker scenarios’ against which should be protected, 3) the utility of the network data in terms of a) which structural properties should be preserved and b) performance on various downstream tasks for which the network will be used, such as detecting communities, and 4) computational cost. The most commonly used methods for the protection of network data have initially been introduced in the context of Statistical Disclosure Control (SDC) [17, 18], aiming to protect tabular data, i.e., data where each entity is described by a number of attributes. In this field one distinguishes between key attributes, quasi-identifiers and sensitive attributes. Key attributes, such as name, personal identifier or phone number, can uniquely identify a person and must be removed from anonymized data. Quasi-identifiers are attributes which may be known to the attacker, who aims to de-anonymize the network, and could possibly be used for identification. Sensitive attributes, such as income or illness, are unknown and of interest to the attacker. Methods for SDC aim to protect against identity disclosure, i.e., finding the person in the tabular dataset, or attribute disclosure, i.e., deriving sensitive attributes of individuals. The utility of the resulting datasets is then measured in terms of general utility, the distribution of attribute values in the data, hence describing the data itself, and specific utility, concerning results of analyzing the data [19].
@@
The most commonly used approaches for anonymizing tabular data are differential privacy [20, 21, 22] and $k$-anonymity [23]. In differential privacy, a mechanism gives noisy answers to user queries such that the privacy of entities is preserved. For $k$-anonymity, the aim is to create a version of the data such that for each entity there are at least $k-1$ equivalent candidates so that any individual can be identified with a probability of at most $1/k$, thereby protecting against identity disclosure. Typically, a value of $k>1$ is used. $k$-Anonymity is then achieved by a combination of generalization by making information on, for example, age or zip code less specific, and suppressing values in the dataset. However, in a $k$-anonymous dataset it could still be possible for an attacker to obtain information if all $k$ entities have the same value for a sensitive attribute or if the distribution in a certain class deviates from the distribution over the entire dataset. Hence, $k$-anonymity in itself does not sufficiently protect against attribute disclosure. This is counteracted by extending $k$-anonymity with $\ell$-diversity [24], $t$-closeness [25] or ($\alpha$, $k$)-anonymity [26]. Other approaches incorporate synthetic data [27] by either generating a new dataset from scratch or combining synthetic data with the actual data. Yet another approach is presented in [28], in which the quasi-identifiers and sensitive attributes are shared in separate tables such that it can not be inferred which set of sensitive attributes corresponds to which set of quasi-identifiers.
Now focusing specifically on network data, based on the methods in SDC several generically applicable approaches have been introduced based on differential privacy [29, 30, 16] as well as $k$-anonymity [31, 32, 11, 33]. Additionally, other network-specific anonymization techniques have been introduced in the literature such as randomization [34, 35, 36] and clustering [37, 32, 38, 39, 40, 41]. These methods ensure that a possible attacker is unable to either identify nodes by ensuring node privacy, or the fact that there is a connection between two entities, i.e., edge privacy.
The approach we focus on in this paper is $k$-anonymity, as this ultimately enables one to publish or share an anonymized version of the network data that can be used for any type of analysis. For $k$-anonymity, one needs to choose both an anonymity measure, and anonymization algorithm. The anonymity measure, which determines when two nodes are equivalent, is used to assess whether the graph satisfies $k$-anonymity. A network is $k$-anonymous when each node is equivalent to at least $k-1$ other nodes. The measure essentially determines the attacker scenario against which a $k$-anonymous network is protected. At first glance, it seems desirable to choose a strict measure, as this protects against more attacker scenarios. However, the stricter the measure, the more properties should be satisfied for two nodes to be equivalent, which means more data utility has to be sacrificed to make nodes anonymous. However, if a too lenient measure is used, hence protecting against a weaker attacker scenario, one takes the risk of underestimating the attacker knowledge and the attacker may still be able to identify entities in the shared network and obtain sensitive information. Usually, real-world networks do not satisfy $k$-anonymity and an anonymization algorithm should be used to perturb the graph and achieve a $k$-anonymous network which can safely be published. This makes the $k$-anonymity approach very flexible. Since each anonymity measure essentially models a specific level of attacker knowledge, it allows one to choose the attacker scenario to protect against accordingly. The anonymization algorithm can be designed to preserve certain network properties considered relevant by the user, such as clustering or centrality of nodes [42]. Utility of the resulting anonymized network can be measured based on how well structural network properties are preserved, such as the clustering coefficient, and performance on various downstream tasks such as the community detection, or finding the most central nodes. While both the anonymity measure and algorithm play an important role, in this paper we focus on the effect of using different anonymization measures.
@@
It is clear that the choice of anonymity measure is of high importance. Although many existing works aim to achieve a $k$-anonymous network for various measures [31, 32, 11, 33] and various surveys have been published on network anonymization in general [43, 44, 45, 15, 46, 47], this paper offers the first systematic comparison of measures for $k$-anonymity. As a key part of this paper, we give an overview of different measures, and compare them both theoretically, and empirically on a wide range of empirical networks. We characterize the anonymity measures based on how far they ‘reach’, i.e., up to where they take structural information into account, and how complete the structural information is and order the measures using a formalized notion of strictness. We empirically compare the measures based on the effectiveness of the attacker scenario considered, the anonymity vs. utility trade-off and runtime. We find that measures with a larger reach, even with minimal structural information, perceive a higher risk of disclosure compared to measures with perfect structural knowledge of the direct neighborhood of a node.
@@
  • We systematically summarize and categorize existing measures for $k$-anonymity in networks (Section 4).
@@
  • We theoretically compare the measures based on reach and the completeness of the structural information they take into account, and derive a formal strictness-based ordering of measures (Section 5).
@@
  • We empirically compare these measures in terms of 1) measured anonymity, 2) the achieved trade-off between anonymity and utility during anonymization, and 3) runtime on a wide range of real-world network datasets (Section 6).
The remainder of the paper is structured as follows. In Section 2, we give an overview of methods introduced in literature to share network data in a privacy aware manner. Then, in Section 3, we provide preliminary concepts and notation, followed by an overview of the considered measures for $k$-anonymity in Section 4. We compare a diverse subset of representative measures theoretically in Section 5, after which we compare these measures empirically on a wide range of network datasets in terms of measured anonymity, utility and runtime in Section 6. Lastly, we conclude the paper and suggest possible future work in Section 7.
2 Related work
In this section, we summarize approaches introduced in literature to share network data while accounting for privacy of individuals represented. We categorize the approaches based on the type of output they produce, being: 1) an interactive mechanism, 2) synthetic data, 3) an intermediate representation, and 4) a perturbed version of the data, which includes $k$-anonymity. We discuss methods in each of these four categories, focusing on the type of output, which attacker scenario one protects against and how utility comes into play. A summary of the methods, corresponding category and type of privacy preserved can be found in Table 1.
Table 1: Categorization of methods for privacy sensitive sharing of networks. $\checkmark$ indicates the method is in this category, $\lozenge$ that it could be in this category, but is less commonly studied as such.
@@ 2 Related work
Differential privacy: queries [30, 48, 49] $\checkmark$ $\lozenge$ $\checkmark$
@@
Clustering [37, 38, 39] $\checkmark$ $\checkmark$
@@
2.1 Interactive mechanism
We first discuss interactive approaches which consists of differential privacy: queries. Approaches based on differential privacy involving synthetic data are discussed in Section 2.2. Differential privacy allows users to ask queries about the network dataset [22]. The user receives an answer to which noise is added, such that it preserves either node or edge privacy. The scenario assumed is that the attacker has knowledge about the entire network except for one node or edge. Based on the answer, it should not be possible to infer whether a specific edge exists in the dataset, when protecting edge privacy, or whether a specific node exists, when protecting node privacy. The amount of noise added is based on the sensitivity of the posed query and ensures privacy is guaranteed. Most works on differential privacy focus on edge privacy [16], and some on node privacy [49, 52]. As the (non)existence of a node can potentially have more effect on the query than the (non)existence of an edge, it is overall more difficult to ensure node privacy. Other works aim to answer a specific user query with as little noise as possible while still achieving the privacy guarantee [30, 48, 49, 15].
A disadvantage of this approach is that a user can only obtain answers to queries that have been implemented, which limits the possible analysis that can be done. Furthermore, when more queries are posed, more noise should be added to ensure differential privacy [15]. This strongly reduces the quality of the answers and with that the utility of the resulting dataset.
2.2 Synthetic data
The second category of approaches generates a synthetic network based on structural properties of the original network. Various models exist to generate networks that capture certain (but not all) real-world network properties. Examples are the Barabási–Albert model [53], which captures the powerlaw degree distribution, or the Watts-Strogatz model [54] which captures the small world property. An example of such a model specifically for achieving privacy, is the dK-graph model, for which the degree of information taken into account can be set with parameter $d$ [55]. For $d=1$, the model uses the degree distribution, and for $d=2$ the joint degree distribution, which incorporates for each edge the degree of the two nodes it connects and how frequently this combination of degrees occurs. However, it was found that using this model alone may not guarantee privacy as many nodes can still be identified in the resulting dK-graph [56].
Work using differential privacy, generates differentially private output which is used to generate synthetic networks. Examples are the dK-graph which uses the joint degree distribution [29, 16], or the exponential random graph model [57]. For this approach, the quality of the resulting network depends both on the quality of the answer provided by the differential privacy mechanism and the graph model used [15]. Additionally, with the increasing use of differential privacy in deep learning [58], recent work proposed a generative deep learning model to generate a network while ensuring edge privacy [50].